OpenStudy (thecuteone25):

MEDAL to anyone who helps the most what is 7/25 as a percent

OpenStudy (fedorable):

do you want an explanation?

OpenStudy (anonymous):

.28

OpenStudy (fedorable):

@Juliusmattos12 don't just give out answers

OpenStudy (thecuteone25):

yes I want an explanation so I know how to do it myself

OpenStudy (fedorable):

ok

OpenStudy (igreen):

Percentages are parts of 100. Make the denominator in \(\sf\dfrac{7}{25}\) over 100.

OpenStudy (igreen):

\(\sf 100 \div 25 =~?\)

OpenStudy (igreen):

@Juliusmattos12 Don't give answers.

OpenStudy (anonymous):

-______- I just started yesterday and when I ask a question everyone gives me the answer

OpenStudy (fedorable):

NEVERMIND THEN

OpenStudy (thecuteone25):

4

OpenStudy (igreen):

Correct, so we multiply both the numerator and denominator by 4. \(\sf 7 \times 4 =~?\) \(\sf 25 \times 4 =~?\)

OpenStudy (igreen):

@Juliusmattos12 That doesn't make it right.

OpenStudy (thecuteone25):

28 and 100

OpenStudy (anonymous):

well I'm sorry I'm saying that cuz I didn't know it was wrong

OpenStudy (igreen):

Correct, so we have: \(\sf\dfrac{28}{100}\) Percentages are parts of 100, so we have \(\sf 28\%\).

OpenStudy (thecuteone25):

thank you

OpenStudy (igreen):

@Juliusmattos12 You were supposed to read the Code of Conduct when you joined. That is stated in there. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

oh

OpenStudy (thecuteone25):

thanks for helping guys

OpenStudy (igreen):

No problem

OpenStudy (anonymous):

lol I didn't help wat u mean "guys" 1 person helped u :/

OpenStudy (thecuteone25):

I was being sarcastic I know only one person helped

OpenStudy (anonymous):

oh